# Robots that Evolve with Us: Modular Co-Design for Personalization, Adaptability, and Sustainability

### v1

@@ 3.2. Workshop Set-up
Introduction. At the beginning, we introduced the idea of modular robots to the participants. We explained that a single robot could change over time by swapping interchangeable parts, such as a head (H), arms (A), an upper body (UB), and a lower body (LB). This meant that instead of imagining different robots for different ages, participants could envision one robot companion that adapts and evolves alongside them throughout their lives. To start, we provided a text-only list of potential modules and functions (e.g., expression generation, climbing mobility, energy harvesting) without showing images (see Table 2, Table 3, Table 4, Table 5). After the participants made their choices, we revealed the visual representations. This sequence encouraged them to focus first on personal needs and functions, before being influenced by appearances. All modules were optional, allowing participants to design either simple or complex configurations. We used the module design from Chen et al. (Chen et al., 2025b) as a basis for the participants to assemble or sketch their designed robot.
Hands-on design. Next, participants considered three major life stages, childhood, adulthood, and older adulthood, and speculated on how a domestic robot might support them at each stage. To emphasize continuity, they were asked to imagine how the same robot could gradually change its form by replacing or adding modules across life stages. For each life stage, they selected relevant functions, retrieved corresponding stickers of modules, and assembled their envisioned robots using stickers on a template (see Figure 1, Figure 2, Figure 3, Figure 4). If the provided modules did not fit their intentions, they were encouraged to sketch their own, fostering creativity and revealing unmet design needs.
@@ 3.3. Data Collection and Analysis
We adopted an interpretive qualitative approach (Adams et al., 2008; Boehner et al., 2007), integrating transcripts, artifacts, and annotations into a single analytic corpus. Two researchers first reviewed a subset of data to generate initial coding categories around robot functions, embodiment choices, and life-stage expectations. We then engaged in iterative open coding and constant comparison, clustering codes into broader categories that captured both convergent patterns and divergent perspectives. Building on this, we applied thematic analysis (Clarke and Braun, 2014; Levinson et al., 2024), a widely used method within HCI’s qualitative research tradition (Fan et al., 2025; Xiao et al., 2025a; Cooper et al., 2022; Xiao et al., 2025b), to identify underlying patterns across the dataset. Through multiple rounds of team discussions, we refined these categories into three overarching themes, which capture not only functional expectations and design preferences but also participants’ values and concerns about long-term human–robot relationships. To ensure credibility, we revisited raw data to ground emerging themes and documented our interpretive decisions through analytic memos.
@@ 4. Findings
Displayed in Figure 8, 9 and 10, we present a total of 65 sketches produced across the two workshops: 45 sketches from 15 participants in Workshop 1, and 20 sketches from 8 participants in Workshop 2. It was noted that P22 preferred a stable look and therefore kept the same design across childhood, adulthood, and older adulthood. P23 only completed the older adulthood stage due to time constraints.
@@ 4.1. Personalization through Modular Design
Modular design allowed participants to treat robots as creative canvases, enabling deep personalization. Participants framed this as essential for maintaining long-term relevance and personal connection across life stage. Participants could selectively choose, skip, remix, or even invent their own modules to construct robots that reflected their individual priorities. These priorities reflect a desire for robots that can grow and transform alongside their users, without losing their core identity.
@@ 4.1.2. Choice and Reconfiguration
A central way modularity supported personalization in our study was by granting participants freedom in both choice and reconfiguration. Rather than being asked to build a “correct” or “complete” robot, participants were encouraged to treat modules as open design materials. This speculative framing gave them permission to question default assumptions about robot form and to imagine alternatives that reflected their own priorities.
@@
Restricting through omission. In contrast, some participants personalized robots by deliberately omitting modules. This approach highlighted different philosophies of support, particularly in childhood designs. For instance, P9 imagined a childhood robot that could help satisfy curiosity without getting “hands dirty,” using A2 arms to fetch objects. Yet P12 resisted such assistance, avoiding arms altogether: “I chose not to add soft arms because I believe children should have the opportunity to explore the world independently. If the robot has arms, it feels as though the robot is doing things for the child, not with a child.” He instead selected H5 (3D projection head) for interactive visuals and UB3 (fabric body) for protection. By leaving out modules, participants could express values of independence or restraint, showing that personalization was not always about “more,” but about aligning designs with personal beliefs about development and care.
@@ 4.2.2. Adaptability for Lifespan Evolution
Adaptability for lifespan evolution thus captured a temporal quality distinct from personalization. Whereas personalization highlighted how different people customized robots in unique ways, adaptability described how one person could sustain a relationship with a single evolving robot across decades. Whether through radical reinvention or selective continuity, participants envisioned modular robots as companions that grow with them, preserving familiarity while transforming to meet changing priorities. This capacity for continuity and reinvention highlights adaptability’s significance: it not only supports long-term usability and ecological sustainability, but also deepens emotional attachment by enabling robots to become lifelong partners rather than disposable machines.
@@ 5.1. Mutual Support through Modular Design: Personalization, Adaptability and Sustainability
The only design-theoretical attempt to integrate these strands so far is the recent Sustainability, Adaptability, and Modularity (SAM) framework (Chen et al., 2025b). SAM usefully highlights the interdependence of these three qualities, positioning sustainability as the overarching aim, modularity as the enabling mechanism, and adaptability as the dynamic capacity to evolve with users’ needs. However, SAM remains a speculative and conceptual vision: it provides a vocabulary for thinking about lifespan-oriented design but does not offer methods, practices, or empirical grounding for how such integration might unfold in real design contexts.
Our contribution lies in extending SAM along three dimensions. First, we recognize SAM as the only design-theoretical framework we could find that explicitly interlinks sustainability, adaptability, and modularity, and we reaffirm its value for lifespan-oriented HRI. Second, we go beyond SAM’s speculative stance by advancing a theoretical extension: we show that modularity can function not only as a technical enabler but as a socio-technical bridge, connecting personalization, adaptability, and sustainability when users are positioned as co-designers. Third, we provide the first empirical grounding of SAM through modular co-design workshops, demonstrating how users themselves assemble, reconfigure, and sustain robots in ways that actualize the framework’s principles. These contributions move SAM from a conceptual vocabulary to an actionable design orientation, showing how lifespan-oriented robot design can be realized in practice.
@@
Personalization as the foundation. We found that personalization provides the entry point for modular engagement. By choosing, skipping, recombining, or inventing modules, participants created robots that embodied their own values, preferences, and priorities. This goes beyond prior HRI research that has focused on tailoring robot roles for specific demographics (e.g., children, older adults) (Klamer and Allouch, 2010; Kory-Westlund and Breazeal, 2019). Instead, personalization in our study highlighted the agency of individuals to define what a robot should be, even within the same demographic group. Such active shaping established a sense of ownership and attachment, laying the emotional groundwork for later practices of repair rather than disposal.
@@
Sustainability as the outcome. Sustainability emerged where personalization and adaptability intersected. Materially, modularity enabled repair, reuse, and incremental upgrading, reducing e-waste and extending lifecycles. Emotionally, continuity and accumulated memories motivated participants to preserve robots rather than replace them. Unlike sustainability research in HRI that has often emphasized ecological tasks performed by robots (e.g., prompting users to recycle)(Scheutz et al., 2021), our findings highlight how the design of the robot itself can embody sustainable principles by fostering both ecological responsibility and emotional continuity. This demonstrates that sustainable HCI requires integrating material and emotional dimensions rather than treating them separately.
@@ 5.2. Design Principles for Future Sustainable HCI/HRI
Provide Open Platforms for Co-Creation. Designers and researchers should move beyond closed, one-size-fits-all systems and instead create open modular platforms that invite users to actively participate in construction and evolution. In our workshops, participants extended the provided modules with new sketches, hybrid designs, and symbolic features, treating modules as open-ended design materials rather than fixed parts (Section 5.1.2). This finding highlights the need for platforms that empower users to experiment, remix, and evolve robots over time. Such openness not only fosters agency and stronger identification with robots but also situates users as co-creators rather than passive consumers, encouraging deeper and more enduring engagement (Desmet and Hekkert, 2007).
Design for Adaptability Within and Across Life Stages. Adaptability should be treated as more than technical flexibility; it is the lived capacity of robots to evolve with the same user’s shifting demands. Our findings showed that users often required different roles within a single stage of life, balancing play and safety in childhood, productivity and care in adulthood, or independence and companionship in older adulthood. Designers should therefore create modules that allow dynamic switching between roles within one life stage, while also enabling smooth transitions across stages. This principle ensures that robots are not frozen in one role but can accompany people through multiple transitions, reflecting the realities of complex and evolving lives.
Foster Sustainability as Both Material and Emotional Practice. Sustainability in modular robots requires addressing both material longevity and emotional continuity. Materially, designs should facilitate easy repair, repurposing, and upgrading of individual modules to prevent full-system replacement and reduce e-waste. Emotionally, continuity can be supported by allowing certain modules or forms to persist across upgrades, enabling users to carry memories and identity connections into the future. As participants noted, familiarity with old modules can evoke nostalgia and strengthen attachment, making robots feel like companions rather than disposable tools. Designers and researchers should therefore enhance the perceived value of robots by encouraging practices of repair and retention, creating products that grow more meaningful over time rather than depreciating with obsolescence (Chapman, 2009; Hebrok, 2014).
@@ 5.3. Limitations and Future Directions
Although childhood was included as a design scenario, children themselves were not directly involved in our workshops. As a result, insights into childhood needs remain speculative, based on adult projections rather than children’s lived experiences. This limits the ecological validity of findings about childhood robot functions. Given that early encounters with robots may shape long-term attitudes, attachment, and sustainability practices, future work should directly involve children to capture their perspectives on play, learning, and safety in robot design.
Our workshops primarily centered on functional roles of modular robots, yet several participants expressed interest in aesthetics and embodiment, such as human-like or culturally resonant appearances. This points to a broader gap in HCI and HRI: while much work has addressed system-level interactions, comparatively little attention has been paid to the design of individual components and their integration into coherent, meaningful forms (DiSalvo et al., 2002). Future research should therefore explore not only how heads, arms, torsos, and bases function, but also how their textures, shapes, and styles align with users’ aesthetic values, cultural expectations, and emotional identities.
Moving forward, future scholars could extend this work by broadening recruitment to include children, culturally diverse participants, and possibly multi-generational families to better capture the inter-generational dynamics of modular robot design. We also see opportunities to iteratively prototype and test modular components that reflect both functional adaptability and aesthetic integration, thereby bridging technical feasibility with cultural and emotional resonance. By shifting attention toward both components and systems, future work can also better realize the promise of modular robotics as a sustainable, adaptable, and inclusive approach to long-term human–robot relationships.

### v2

@@ 3.2. Workshop Set-up
Introduction. At the beginning, we introduced the idea of modular robots to the participants. We used the module design set and its associated tables of potential functions from Chen et al. (Chen et al., 2025b) as the shared design basis for the participants to assemble or sketch their designed robot. Heads (H), arms (A), an upper body (UB), and a lower body (LB) served as a basic structural scaffold, but participants were not required to use all of them. To keep the focus on personal needs and functions rather than appearance, participants first selected desired functions from text-only tables (e.g., expression generation, climbing mobility, energy harvesting) without showing images (see Table 2, Table 3, Table 4, Table 5). Only after they had made these choices did we reveal the corresponding visual module designs and invite them to assemble or sketch their robots. This sequence encouraged participants to imagine one robot that could be reconfigured to support different life stages, rather than designing separate robots for each age.
Hands-on design. Next, participants considered three major life stages, childhood, adulthood, and older adulthood, and were asked to design a modular robot for themselves. They imagined how the same robot might assist and accompany themselves rather than designing for an abstract or generic user. In doing so, participants chose whether to speak from the vantage point of their younger self, their current adult self, or their anticipated older self. They speculated on how a modular robot might support their everyday lives across these stages. We deliberately framed the task around home and nearby everyday settings rather than industrial or large-scale public environments. Within this framing, participants were free to assign any roles they found meaningful, ranging from tool-like helpers and task-oriented assistants to more social or emotionally supportive partners. To emphasize continuity, they were asked to imagine how the same robot could gradually change its form by replacing or adding modules across life stages. For each life stage, they selected relevant functions, retrieved corresponding stickers of modules, and assembled their envisioned robots using stickers on a template (see Figure 1, Figure 2, Figure 3, Figure 4). If the provided modules did not fit their intentions, they were encouraged to sketch their own, fostering creativity and revealing unmet design needs.
@@ 3.3. Data Collection and Analysis
We adopted an interpretive qualitative approach (Adams et al., 2008; Boehner et al., 2007), integrating transcripts, artifacts, and annotations into a single analytic corpus. Two researchers first reviewed a subset of data to generate initial coding categories around robot functions, embodiment choices, and life-stage expectations. We then engaged in iterative open coding and constant comparison, clustering codes into broader categories that captured both convergent patterns and divergent perspectives. Building on this, we applied thematic analysis (Clarke and Braun, 2014; Levinson et al., 2024), a widely used method within HCI’s qualitative research tradition (Fan et al., 2025; Cooper et al., 2022; Xiao et al., 2025), to identify underlying patterns across the dataset. Through multiple rounds of team discussions, we refined these categories into three overarching themes, which capture not only functional expectations and design preferences but also participants’ values and concerns about long-term human–robot relationships. To ensure credibility, we revisited raw data to ground emerging themes and documented our interpretive decisions through analytic memos.
@@ 4. Findings
We present a total of 65 sketches produced across the two workshops: 45 sketches from 15 participants in Workshop 1, and 20 sketches from 8 participants in Workshop 2. The complete collection of sketches is shown in Appendix A. In the main text, we focus on selected, annotated examples to illustrate how participants used modular configurations to realize personalization, adaptability, and sustainability. It was noted that P22 preferred a stable look and therefore kept the same design across childhood, adulthood, and older adulthood. P23 only completed the older adulthood stage due to time constraints.
@@ 4.1. Personalization through Modular Design
Modular design allowed participants to treat robots as creative canvases, enabling deep personalization. Participants framed this as essential for maintaining long-term relevance and personal connection across life stage. Rather than simply “liking customization,” they used modularity to negotiate whether a robot still felt appropriately matched to their changing roles and self-image at each life stage. Participants could selectively choose, skip, remix, or even invent their own modules to construct robots that reflected their individual priorities. These priorities reflect a desire for robots that can grow and transform alongside their users, without losing their core identity.
@@ 4.1.2. Choice and Reconfiguration
A central way modularity supported personalization in our study was by granting participants freedom in both choice and reconfiguration. Rather than being asked to build a “correct” or “complete” robot, participants were encouraged to treat modules as open design materials. This speculative framing gave them permission to question default assumptions about robot form and to imagine alternatives that reflected their own priorities. Figure 9 shows examples of designs with redundant parts, certain bodyparts omitted, or reconfigurations of parts for other purposes.
@@
Restricting through omission. In contrast, some participants personalized robots by deliberately omitting modules. This approach highlighted different philosophies of support, particularly in childhood designs. For instance, P9 imagined a childhood robot that could help satisfy curiosity without getting “hands dirty,” using A2 arms to fetch objects. Yet P12 resisted such assistance, avoiding arms altogether: “I chose not to add soft arms because I believe children should have the opportunity to explore the world independently. If the robot has arms, it feels as though the robot is doing things for the child, not with a child.” He carried this stance into adulthood, again omitting arms. He configured H4 (sensor head), UB2 (storage body), and LB4 (wheeled base) to prioritize awareness and mobility over manipulation, arguing that the robot should support independence rather than perform tasks on his behalf. By leaving out modules, participants could express values of independence or restraint, showing that personalization was not always about “more,” but about aligning designs with personal beliefs about development and care.
@@ 4.2.2. Adaptability for Lifespan Evolution
Adaptability for lifespan evolution thus captured a temporal quality distinct from personalization, whereas personalization highlighted how different people customized robots in unique ways, adaptability described how one person could sustain a relationship with a single evolving robot across decades. Whether through radical reinvention or selective continuity, participants envisioned modular robots as companions that grow with them, preserving familiarity while transforming to meet changing priorities. This capacity for continuity and reinvention highlights adaptability’s significance: it not only supports long-term usability and ecological sustainability, but also deepens emotional attachment by enabling robots to become lifelong partners rather than disposable machines.
@@ 5.1. Mutual Support through Modular Design: Personalization, Adaptability and Sustainability
The only design-theoretical attempt to integrate these strands so far is the recent Sustainability, Adaptability, and Modularity (SAM) framework (Chen et al., 2025b). SAM usefully highlights the interdependence of these three qualities, positioning sustainability as the overarching aim, modularity as the enabling mechanism, and adaptability as the dynamic capacity to evolve with users’ needs. Our work takes this agenda as a starting point and asks how it plays out when people actually co-design modular robots for their own lives.
Our contribution lies in complementing SAM along three dimensions. First, we recognize SAM as a design-theoretical framework that explicitly interlinks sustainability, adaptability, and modularity, and we reaffirm its value for lifespan-oriented HRI. At the same time, our analysis highlights a specific limitation: SAM articulates sustainability and adaptability as desirable system properties, but does not make clear for whom or how these properties are oriented toward particular life courses. Second, we address this limitation by introducing personalization as a human-centered process that orients SAM’s system-level aims toward specific people and life situations. In our data, modular configurations were consistently tied to participants’ concrete roles, relationships, bodies, and constraints. We use personalization to describe how a robot’s changing form and behavior are continually adjusted to these situated values and priorities. Our findings suggest that, in lifespan-oriented HRI, the connection between adaptability and sustainability is strongly mediated by meaning-making—how people come to see a robot as “for me,” “part of my life,” and “worth keeping.” When modular reconfigurations express identity, roles, and values, they cultivate ownership and attachment, which in turn increases people’s willingness to maintain, repair, and carry modules forward rather than abandon the robot. We therefore treat personalization not as an add-on dimension, but as an important human-centered mechanism that gives direction to adaptability and motivates sustained use. Without this lens, adaptability remains directionless and sustainability risks being understood only in technical terms (e.g., longer lifetimes) rather than as something that is worth maintaining and repairing. In PAS, modularity therefore functions not only as a technical enabler but as a socio-technical bridge that connects personalization, adaptability, and sustainability when people are positioned as co-designers. Third, we provide the first empirical grounding of SAM through modular co-design workshops, demonstrating how people themselves assemble, reconfigure, and sustain robots in ways that actualize the framework’s principles. These contributions move SAM from a conceptual vocabulary to an actionable design orientation, and distinguish our PAS framework as a human-centered, empirically grounded account of how lifespan-oriented robot design can be realized in practice.
@@
Personalization as the foundation. We found that personalization emerged as the entry point for modular engagement. By choosing, skipping, recombining, or inventing modules, participants created robots that embodied their own values, preferences, and priorities. This goes beyond prior HRI research that has focused on tailoring robot roles for specific demographics (e.g., children, older adults) (Klamer and Allouch, 2010; Kory-Westlund and Breazeal, 2019). Instead, personalization in our study highlighted the agency of individuals to define what a robot should be, even within the same demographic group. Such active shaping established a sense of ownership and attachment, laying the emotional groundwork for later practices of repair rather than disposal.
@@
Sustainability as the outcome. Sustainability was not introduced to participants as an explicit design goal. Instead, it surfaced in how they described keeping, repairing, and repurposing a single robot over time. Materially, we understand that modularity enables repair, reuse, and incremental upgrading, reducing e-waste and extending lifecycles. Emotionally, the participants framed continuity, memories, and attachment as reasons to preserve a familiar robot instead of replacing it with a newer model. In their reflections, participants often described these decisions in terms of making one robot last, keeping a trusted companion, and avoiding unnecessary replacement. We therefore interpret these practices as a form of sustainability, reading them through a sustainable HCI lens. Unlike sustainability research in HRI that has often emphasized ecological tasks performed by robots (e.g., prompting people to recycle) (Scheutz et al., 2021), our findings highlight how the design of the robot itself can embody sustainable principles by fostering both ecological responsibility and emotional continuity. This perspective resonates with work in HRI that examines robots as mediators of destruction, catharsis, and emotional release rather than purely utilitarian tools (Luria et al., 2020). This suggests that long-lived companions must be able to accommodate intense, sometimes messy interactions without being treated as disposable. This demonstrates that sustainable HCI requires integrating material and emotional dimensions rather than treating them separately.
@@ 5.2. Design Principles for Future Sustainable HCI/HRI
Provide Open Platforms for Co-Creation. Designers and researchers should move beyond closed, one-size-fits-all systems and instead create open modular platforms that invite users to actively participate in construction and evolution. In our workshops, participants extended the provided modules with new sketches, hybrid designs, and symbolic features, treating modules as open-ended design materials rather than fixed parts (Section 5.1.2). This finding highlights the need for platforms that empower users to experiment, remix, and evolve robots over time. Such openness not only fosters agency and stronger identification with robots but also situates users as co-creators rather than passive consumers, encouraging deeper and more enduring engagement (Desmet and Hekkert, 2007), (Alves-Oliveira et al., 2021; Cui et al., 2024).
Design for Adaptability Within and Across Life Stages. Adaptability should be treated as more than technical flexibility; it is the lived capacity of robots to evolve with the same user’s shifting demands in both function and form. Our findings showed that users often required different roles within a single stage of life, balancing play and safety in childhood, productivity and care in adulthood, or independence and companionship in older adulthood. At the same time, adaptability should guard against robots that look and behave exactly the same for decades. Companions that never change can begin to feel oddly out of sync with people’s lives, raising a sense of “social uncanniness” when they do not age or transform alongside their partners (Hoffman, 2020). Designing modules that can be added, removed, or visually refreshed over time allows robots to register life transitions—subtle shifts in appearance, capabilities, or interaction style that mirror peoples’ own changing identities. Designers should therefore create modules that allow dynamic switching between roles within one life stage, while also enabling smooth transitions across stages. This principle ensures that robots are not frozen in one role but can accompany people through multiple transitions, reflecting the realities of complex and evolving lives.
Foster Sustainability as Both Material and Emotional Practice. Sustainability in modular robots requires addressing both material longevity and emotional continuity. Materially, designs should facilitate easy repair, repurposing, and upgrading of individual modules to prevent full-system replacement and reduce e-waste. Emotionally, continuity can be supported by allowing certain modules or forms to persist across upgrades, enabling users to carry memories and identity connections into the future. As participants noted, familiarity with old modules can evoke nostalgia and strengthen attachment, making robots feel like companions rather than disposable tools. Evidence from care settings with the therapeutic robot PARO similarly shows that long-term engagement with a single, familiar robot can help reduce agitation and loneliness and support more positive care experiences for older adults (Šabanović et al., 2013b; Hung et al., 2019). Designers and researchers should therefore enhance the perceived value of robots by encouraging practices of repair and retention, creating products that grow more meaningful over time rather than depreciating with obsolescence (Chapman, 2009; Hebrok, 2014).
@@ 5.3. Limitations and Future Directions
Although childhood was included as a design scenario, children themselves did not participate in our workshops, since we believe that understanding adult and older adult perspectives first can better inform a future study that includes children as participants. All childhood designs were therefore based on adult projections rather than children’s lived experiences. This constrains the ecological validity of our findings for early life stages, especially around play, learning, and safety. We began with this adult-centred, lifespan sample as an experimental first step in developing the PAS framework: adults can speak both as former children and as current or prospective caregivers, which allowed us to articulate and stress-test the dimensions of personalization, adaptability, and sustainability within a single, coherent lifespan narrative. At the same time, this vantage point means that our accounts of early life stages should be read as provisional and situated in adult imaginaries, rather than as direct evidence of children’s perspectives. Future work should complement this view by directly involving children and teenagers—for example, through short, age-appropriate co-design activities or probes—to triangulate and extend our findings about how modular robots might support early-life play, learning, and safety.
Our workshops primarily centered on the functional roles of modular robots, yet several participants expressed interest in aesthetics and embodiment, such as human-like or culturally resonant appearances. This points to a broader gap in HCI and HRI—while much work has addressed system-level interactions, comparatively little attention has been paid to the design of individual components and their integration into coherent, meaningful forms (DiSalvo et al., 2002). At the same time, our contribution is grounded in conceptual and graphical co-design probes rather than a tangible modular prototype. Participants imagined how a single robot might adapt with them over decades, but neither embodiment nor technical feasibility was tested. As a result, our claims about lifelong adaptability and modular growth should be understood as design-oriented and hypothetical rather than as evidence of a working system. To substantiate these arguments, future research should include material explorations and iterative prototyping—even at low fidelity—that examine how specific module designs, connectors, textures, and forms perform in real use, and how they align with peoples’ aesthetic values, cultural expectations, and emotional identities. Such work would help clarify which aspects of personalization, adaptability, and sustainability are practically achievable, and what trade-offs arise when modular robots are implemented in concrete contexts.
Moving forward, future research should also reconsider the structural assumptions built into the modules themselves, and further relax or even abandon the humanoid scaffold that underpins our current module set. In this study, we deliberately started from a body-like construction grammar precisely to surface how participants reproduce or resist familiar expectations about “humanoid” robots. Our findings show that people are willing to bend and break this structure when modularity allows them to embed additional functions and meanings. Building on metaphor-driven approaches that use open-ended prompts to expand robot imaginaries beyond human-like or pet-like archetypes, such as Alves-Oliveira et al.’s collection of metaphors for HRI (Alves-Oliveira et al., 2021), future co-design studies could instead begin from non-anthropomorphic, environment-centered, or distributed module grammars. This would allow us to investigate how lifespan-oriented modularity plays out when the underlying structure itself is no longer tied to a human body form.